OpenStudy (anonymous):

Simplify the expression (2x − 9)(x + 6).

OpenStudy (anonymous):

2x^2+3x-54

OpenStudy (anonymous):

i agree with @TheTerrence

OpenStudy (texaschic101):

You have seen the answer, but do you know how to get it ? Use the FOIL method. FOIL stands for first,outside,inside,last. (2x - 9)(x + 6) First -- multiply 1st terms in each set....2x * x = 2x^2 Outside -- multiply outside terms in each set...2x * 6 = 12x Inside -- multiply inside terms in each set...-9 * x = -9x Last -- multiply last terms in each set...-9 * 6 = -54 Now put it all together.. 2x^2 + 12x - 9x - 54 --- combine like terms 2x^2 + 3x - 54 There you go...just in case you were wondering how to get the answer :)